# EndoChat: Grounded Multimodal Large Language Model for Endoscopic Surgery

### v1

@@ Acknowledgements
Author contributions statement
Guankun Wang: Conceptualization, Methodology, Validation, Investigation, Original draft preparation. Long Bai: Conceptualization, Methodology, Investigation, Original draft preparation. Junyi Wang: Validation, Investigation, Dataset Generation. Kun Yuan: Methodology, Original draft preparation. Tianxu Jiang: Dataset Generation. Xiting He: Manuscript Review. Jinlin Wu: Computing Resource Support, Manuscript Review. Zhen Chen: Manuscript Review. Zhen Li: Endoscopist Evaluation, Manuscript Review. Hongbin Liu: Manuscript Review. Nicolas Padoy: Manuscript Review. Nassir Navab: Manuscript Review. Hongliang Ren: Conceptualization, Supervision, Review and editing, Resources.

### v2

@@ 7 Acknowledgements
8 Competing interests
The authors declare no competing interests.
9 Author contributions statement
Guankun Wang: Conceptualization, Methodology, Validation, Investigation, Writing – original draft. Long Bai: Conceptualization, Methodology, Investigation, Writing – original draft. Junyi Wang: Validation, Investigation, Dataset Generation, Writing – review and editing. Kun Yuan: Methodology, Writing – original draft. Zhen Li: Endoscopist Evaluation, Writing – review and editing. Tianxu Jiang: Dataset Generation, Writing – review and editing. Xiting He: Software, Writing – review and editing. Jinlin Wu: Computing Resource Support, Writing – review and editing. Zhen Chen: Conceptualization, Writing – review and editing. Zhen Lei: Formal analysis, Writing – review and editing. Hongbin Liu: Formal analysis, Writing – review and editing. Jiazheng Wang: Conceptualization, Writing – review and editing. Fan Zhang: Conceptualization, Writing – review and editing. Nicolas Padoy: Data curation, Writing – review and editing. Nassir Navab: Conceptualization, Data curation, Writing – review and editing. Hongliang Ren: Conceptualization, Supervision, Writing – original draft, Resources.